OpenStudy (anonymous):

the equation of the line tangent to y=2secx at ax=pi/4 is ?

OpenStudy (anonymous):

ok so the next step is this y=2sec(pi/4)tan(pi/4), right???

hartnn (hartnn):

you mean \(\large y'|_{x=\pi/4} = 2\sec (\pi/4)\tan (\pi/4)\) then yes, you're correct. and that will give you slope = m

OpenStudy (anonymous):

is the slope 2 ?? cuz thats whats i get

hartnn (hartnn):

not actually, sec pi/2 is sqrt 2

hartnn (hartnn):

***sec pi/4 = sqrt 2

OpenStudy (anonymous):

(2/